Assume a sample of 100 units provide a 95% confidence interval [10,20]. Which of the following is TRUE? Question 1 options: margin of error = 5, sample mean = 15 margin of error = 15, sample mean = 5 margin of error = 10, sample mean = 15 margin of error = 15, sample mean = 30

Question 2 (1 point) Assume a sample of 100 units provide a 95% confidence interval [10,20]. Which of the following could correspond to a 99% confidence interval? Question 2 options: [15,30] [12,18] [10,30] [7,23]

Question 3 (1 point) Which of the following is FALSE if you increase the sample size? Question 3 options: the margin of error should go up , you should get a narrower interval the margin of error should go down , you should get a wider interval the margin of error should go down , you should get a narrower interval the margin of error should go up , you should get a wider interval

Question 4 (1 point) A new system promises to reduce the average delivery time of 22 hours. After the implementation of the new system, a sample of 100 deliveries showed a sample mean of 21.5 hours with a standard deviation of 2.5 hours. Select the CORRECT formulation. Question 4 options: Ho: μ ≥ 22 Ha: μ < 22 Ho: μ < 21.5 Ha: μ ≥ 21.5 Ho: μ < 22 Ha: μ ≥ 22 Ho: μ ≥ 21.5 Ha: μ < 21.5

Question 5 (1 point) A new system promises to reduce the average delivery time of 22 hours. After the implementation of the new system, a sample of 100 deliveries showed a sample mean of 21.5 hours with a standard deviation of 2.5 hours. What statistic would you use to calculate the p-value? What type of test is this? Question 5 options: I use z. This is a two-tailed test I use z. This is a one-tailed test I use t. This is a two-tailed test I use t. This is a one-tailed test
Answer #1

n=100 Interval: X I MOG X-MOE = 10 XAME = 20 - 2X=30 X=15 MOE=20-15=5 . Sample mean 15, MOE=5 99.). c. I would be wider as mo